∫e lnx sinx dx

∫elnxsinx dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int e^{lnx}sinxdx
I = \int xsinxdx
Apply ILATE Rule
I = -xcosx - \int (-cosx)dx
I = -xcosx +\int cosxdx
I = -xcosx +sinx + c
I = (sinx-xcosx) + c